52(x-1)+5x+1=0.04

25


1
Expert's answer
2021-02-16T15:00:45-0500

"5^{2(x-1)}+5^{x+1}=\frac{4}{100}=\frac{1}{5^2}"


"5^{2x}*5^{-2}+5^x*5=5^{-2}"

multiplying both side by "5^{2}" , we get

"5^{2x}+5^x*5^{2+1}=1" ("a^p*a^q=a^{p+q}" )

"5^x*5^x+5^x*5^3=1" ; on dividing the equation by 5x in both side, we get

"5^x+5^3=\frac{1}{5^x}" or "5^x-\frac{1}{5^x}=-125"


let "5^x=y" , then the above equation can be written as,

"y-\frac{1}{y}=-125" or "y^2+125y-1=0"


"y=\frac{-125\pm\sqrt{125^2-4*1*(-1)}}{2*1}" "=\frac{-125\pm(125.015999)}{2}"

y= -125.008 or y= 0.0079995

we know that "y=5^x" , for any value of x, gives the value of y always positive. so

y = 0.0079995 is only the valid value.

so, "y=0.0079995=5^x"

taking log (base 10) both side of equation, we get

"log(0.0079995)=log(5^x)"

"-2.09693716 = x * log(5)" "(log(a^p)=p*log(a))"


"x=\frac{-2.09693716}{0.698970004}"


"x=-3.00003884"